# TAROT: Targeted Data Selection via Optimal Transport

### v1

@@ 5.3 TAROT for Motion Prediction
  • Can targeted selection improve the performance of state-of-the-art motion prediction models on a target domain?
Method Ranking ($\downarrow$) minADE5555 ($\downarrow$)
Wayformer-TAROT 1 0.92
Wayformer-All [32] 2 0.94
MTR-UniTraj [15] 3 0.96
Goal-LBP [45] 4 1.02
SemanticFormer [37] 5 1.14
@@
Target and Candidate Datasets. We use the UniTraj framework [15] for unified training and evaluation across multiple datasets, including Waymo Open Motion (WOMD) [14], Argoverse 2 [41], nuScenes [5], and nuPlan [18].
This experiment focuses on improving model performance on the nuScenes validation set through targeted data selection. The nuScenes training set ($32k$ samples) serves as the target dataset, while the candidate pool comprises WOMD, Argoverse 2, and nuPlan. From nuPlan, we filter trajectories with a moving distance over 2 meters, yielding $1000k$ samples. WOMD and Argoverse 2 use their official training splits, containing $1800k$ and $180k$ samples, respectively. The evaluation is conducted on the nuScenes validation set. This cross-dataset task involves challenges like non-overlapping geographic regions and varying annotation formats. We explore whether TAROT can effectively curate multi-source data to enhance nuScenes validation performance.
Baselines. We compare against the same baselines in Sec. 5.2.
Data Selection and Model Training. We employ AutoBots [16], a lightweight ($1.5M$ parameters) attention-based prediction model, for data selection. Autobots is pretrained on candidate and target datasets for 50 epochs, and gradients from the final checkpoint guide selection.
We evaluate selection ratios of 5%, 20%, 50%, and OTM. OTM selects approximately 31%, 53%, and 22% from WOMD, Argoverse 2, and nuPlan, covering 29.8% of the full dataset. The selected data is used to train the Wayformer [32] model ($15M$ parameters) for 100 epochs. To highlight that TAROT can also achieve performance gains with reduced training data and iterations, the data weighting method (Sec. 4.4) is not applied.
Evaluation Metric. We use the Brier Minimum Final Displacement Error (Brier-minFDE) metric for evaluation, which is used in the Argoverse 2 [41] challenge. We report the best validation performance for each method.
Main Results. Fig. 7 compares the performance of Wayformer trained on datasets with different selection ratios. Key findings include:
  • Effectiveness of TAROT: TAROT consistently outperforms other baselines. Influence-estimation based heuristics, such as LESS and DsDm, perform well at low selection ratios (5%) but lose effectiveness as the ratio increases, emphasizing the benefits of the distribution-matching approach. Additionally, OTM selection outperforms training on the entire dataset and other selection ratios, delivering global optimal results.
@@
  • Improvement on leaderboard performance. Notably, Sec. 5.3 shows that Wayformer trained on TAROT-selected data achieves first place on the nuScenes leaderboard, surpassing MTR-UniTraj [15], which relies on large-scale datasets ($2000k$ samples). These results demonstrate that targeted data selection offers a promising alternative to data scaling for breaking performance boundaries.
@@ 5.4 TAROT for Instruction Tuning
Previous experiments demonstrate that TAROT is particularly effective when dealing with complex target distributions. In this section, we evaluate whether TAROT can also deliver strong performance in large language model (LLM) tasks, where the target distribution is relatively simple and existing methods have already shown high effectiveness.
Target and Candidate Datasets. To ensure a fair comparison and maintain consistency with established benchmarks, we replicate the experimental setup described by Xia et al. [42]. We utilize the same candidate dataset, comprising Flan V2 [30], CoT [40], Dolly [8] and Open Assistant 1 [28], with MMLU [21, 20] and BBH [38] serving as the target tasks for evaluation.
Data Selection and Model Training. We conduct experiments using Llama-3.1-8B [12] and Qwen-2.5-7B [39] as the base models for both data selection and fine-tuning. To evaluate the transferability of our selection method, we also fine-tune a Qwen-2.5-7B using data samples selected by Llama-3.1-8B. Data selected by TAROT-OTM is weighted using OT potentials (Sec. 4.4) with $R=0.5\%N$. More details are provided in the supplementary materials.
@@
Main Results. Tab. 2 represents the performance of TAROT compared with baselines. Among them, LESS requires the subtask labels of target samples, whereas TAROT operates without this requirement. Our key findings include:
  • Effectiveness and transferability of TAROT. TAROT consistently outperforms all baseline methods and models trained on the full dataset without relying on subtask labels. Besides, data samples selected by Llama-3.1-8B enhances performance for Qwen-2.5-7b, indicating strong transferability across different model architectures.
@@ 7 Acknowledgements
Supplementary Material
8 More Implementation Details
8.1 Influence Estimation
ZCA Whitening. As discussed in Sec. 5.1, we use ZCA whitening to achieve higher Local Data Scaling (LDS) scores due to its ability to preserve the data structure. However, our empirical analysis revealed that ZCA whitening has minimal impact on data selection performance. Therefore, for the data selection experiments, we opted for the more computationally efficient Cholesky whitening.
@@ 8.1 Influence Estimation
Linear Data Modeling Score (LDS). As mentioned in Sec. 5.1, we employ LDS as evaluation metric for influence estimation. LDS is computed by the Spearman rank correlation between the model outputs of models trained on randomly selected subsets and the average of attribution scores as predictions of those model outputs. For Cifar-10, we directly reuse the provided outputs of models trained on different subsets in the Github repository of TRAK [34], and utilize these outputs to calculate LDS. For nuScences, we randomly select 100 subsets of the candidate datasets. Each subset is sampled to be 50% of the size of the candidate dataset. Then we independently train 100 models on these subsets and utilize their model outputs to calculate LDS.
@@
8.2 Semantic Segmentation
Tab. 3 shows detailed hyperparameters for the semantic segmentation experiments.
8.3 Motion Prediction
We use the UniTraj [15] framework for motion prediction experiments. Detailed experiment settings are provided in Tab. 4.
8.4 Instruction Tuning
We follow the settings and hyperparameters in LESS [42]. Details are provided in Tab. 5
9 Details for Algorithms
We present the detailed data selection algorithm described in  Sec. 4.3
10 More Experiment Results
10.1 Ablation Study on Whitened Feature Distance
Table 6: Ablation Study of WFD on nuScenes with ensemble size ranging from 1 to 50. The evaluation metric is LDS.
1 10 50
w/o whitened 0.007 0.010 0.017
w/o norm 0.024 0.060 0.097
w/o whitened+norm 0.005 0.008 0.015
WFD 0.028 0.065 0.098
@@ 10.1 Ablation Study on Whitened Feature Distance
In Tab. 6, we conduct ablation experiments by removing key components to explore their impacts on WFD. ’w/o whitened’ means we do not perform whitening to the projected gradients and utilize the original projected gradients to calculate the distance. ‘w/o norm’ means we compute the L2 distance of the gradient vectors without normalization. Results demonstrate that removing the proposed components of our whitened feature distance will lead to a performance drop, demonstrating the effectiveness of our influence estimation design.
10.2 More Qualitative Results of WFD
@@
We provide additional qualitative results of WFD in Figs. 8 and 9. One clear observation is that the most supportive samples tend to closely resemble the target sample in both the input and output spaces. In motion prediction, the most negative samples also share similarities with the target sample in the input space (e.g., map and history trajectories); however, their ground truth trajectories often exhibit minor differences. For instance, in the first row of samples, the most negative examples demonstrate behaviors such as turning left or driving in the right-most lanes, whereas the target sample follows a trajectory in the lane adjacent to the right-most lane. These qualitative findings align with the definition of influence estimation: the most positive samples guide the model to make predictions similar to the target sample, while the negative samples encourage divergent predictions.

### v2

@@ 4.3 TAROT for Motion Prediction
  • Is TAROT generic? Compared to traditional computer vision tasks (e.g., image classification), motion prediction involves more complex inputs (e.g., maps, trajectories) and outputs (e.g., predicted trajectories and confidence scores). Is TAROT still effective in a more challenging setting?
@@
Datasets. We use the UniTraj framework (Feng et al., 2024) for unified training and evaluation across multiple datasets, including Waymo Open Motion (WOMD) (Ettinger et al., 2021), Argoverse 2 (Wilson et al., 2021), nuScenes (Caesar et al., 2020), and nuPlan (H. Caesar, 2021). The nuScenes training set ($32k$ samples) serves as the target dataset, while the candidate pool comprises WOMD, Argoverse 2, and nuPlan. From nuPlan, we filter trajectories with a moving distance over 2 meters, yielding $1000k$ samples. We use the official training splits of WOMD and Argoverse 2, including $2000k$ samples. The evaluation is conducted on the nuScenes validation set. This cross-dataset task involves challenges like non-overlapping geographic regions and varying annotation formats. We explore whether TAROT can effectively curate multi-source data to enhance nuScenes validation performance.
Baselines. We utilize the same baselines in Section 4.2.
Data Selection and Model Training. We employ AutoBots (Girgis et al., 2021), a lightweight ($1.5M$ parameters) attention-based prediction model, for data selection. Autobots is pretrained on candidate and target datasets for 50 epochs, and gradients from the final checkpoint guide selection. We evaluate selection ratios of 5%, 20%, 50%, and OTM. OTM selects approximately 31%, 53%, and 22% from WOMD, Argoverse 2, and nuPlan, covering 29.8% of the full dataset. The selected data is used to train the Wayformer (Nayakanti et al., 2023) model ($15M$ parameters) for 100 epochs. To highlight that TAROT can also achieve performance gains with reduced training data and iterations, the data weighting method (Section 3.4) is not applied.
Evaluation Metric. We use Brier Minimum Final Displacement Error (Brier-minFDE) metric for evaluation, which is used in the Argoverse 2 (Wilson et al., 2021) challenge. We report the best validation performance for each method.
Main Results. Figure 7 compares the performance of Wayformer trained on datasets with different selection ratios. Key findings include:
  • Effectiveness of TAROT: TAROT consistently outperforms other baselines. Influence-estimation based heuristics, such as LESS and DsDm, perform well at low selection ratios (5%) but lose effectiveness as the ratio increases, emphasizing the benefits of the distribution-matching approach. Additionally, OTM selection outperforms training on the entire dataset and other selection ratios, delivering global optimal results. Moreover, TAROT archives the top 1 performance on the nuScenes leaderboard. Results are shown in the Appendix D.
@@ 4.4 TAROT for Instruction Tuning
Previous experiments demonstrate that TAROT is particularly effective when dealing with complex target distributions. In this section, we evaluate whether TAROT can also deliver strong performance in LLM tasks, where the target distribution is relatively simple and existing methods have already shown high effectiveness.
Datasets. To ensure a fair comparison and maintain consistency with established benchmarks, we replicate the experimental setup described by Xia et al. . We utilize the same candidate dataset, comprising Flan V2 (Longpre et al., 2023), CoT (Wei et al., 2022), Dolly (Conover et al., 2023) and Open Assistant 1 (Köpf et al., 2024), with MMLU (Hendrycks et al., 2021b, a) and BBH (Suzgun et al., 2023) serving as the target tasks for evaluation.
Data Selection and Model Training. We conduct experiments using Llama-3.1-8B (Dubey et al., 2024) and Qwen-2.5-7B (Team, 2024) as the base models for both data selection and fine-tuning. To evaluate the transferability of our selection method, we also fine-tune a Qwen-2.5-7B using data samples selected by Llama-3.1-8B. Data selected by TAROT-OTM is weighted using OT potentials (Section 3.4) with $R=0.5\%N$. More details are provided in the supplementary materials.
@@
Main Results. Table 1 represents the performance of TAROT compared with baselines. Among them, LESS requires the subtask labels of target samples, whereas TAROT operates without this requirement. Our key findings include:
  • TAROT is effective, generalizable, and transferable. TAROT consistently outperforms all baseline methods and models trained on the full dataset without relying on subtask labels. Besides, data samples selected by Llama-3.1-8B enhances performance for Qwen-2.5-7b, indicating strong transferability across different model architectures.
@@
4.5 Computational Complexity
The computational complexity of TAROT mainly arises from two aspects: (1) gradient feature computation and (2) data selection. The former shares the same complexity as previous methods that use projected gradients (Park et al., 2023; Liu et al., 2024; Engstrom et al., 2024). The latter introduces additional computational overhead due to computation of OT distance. We compare the wall-clock runtime of TAROT with baselines and provide a detailed asymptotic complexity analysis in the Appendix G.
4.6 Ablation Study
Table 2: Ablation Study of WFD on nuScenes with ensemble size ranging from 1 to 50. Evaluation metric is LDS.
Ensemble size 1 10 50
w/o whitened 0.007 0.010 0.017
w/o norm 0.024 0.060 0.097
w/o whitened+norm 0.005 0.008 0.015
WFD 0.028 0.065 0.098
Figure 8: Ablation study for the choice of k. The figure shows the performance of Wayformer under different k.
Whitened Feature Distance. We conduct an ablation study to analyze the impact of key components on WFD. The results are presented in Table 2. Specifically, w/o whitened indicates that we skip the whitening step and instead use the original projected gradients to compute the distance. w/o norm means that we compute the L2 distance of the gradient vectors without normalization. The results demonstrate that removing these proposed components leads to a performance drop, highlighting the effectiveness of our influence estimation design.
K-Fold Selection in OTM. As introduced in Section 3.3, TAROT employs k-fold cross-validation for OTM, where a larger $k$ results in less alignment with the target distribution. We investigate how the choice of $k$ affects data selection in the motion prediction setting (Section 4.3). Specifically, we perform experiments with OTM selection using $k$ values ranging from 1 to 20, evaluating the model’s performance on both the test set and the target set. The results, shown in Figure 8, indicate that as $k$ increases, the model’s performance on the test set improves, reaching its peak at $k=10$. Meanwhile, performance on the target set steadily declines, suggesting that the selected data becomes less aligned with the target distribution as $k$ increases.
5 Related Work
We review the progress on data selection in this section. More related work is deferred to App. A. Coreset selection aims to identify a dataset subset for optimal in-domain performance (Xia et al., 2022; Mirzasoleiman et al., 2020; Borsos et al., 2020; Killamsetty et al., 2021; Guo et al., 2022; Xiao et al., 2024), assuming validation and candidate data share the same distribution. Our work shifts focus to targeted selection, optimizing performance on a distinct target domain. Prior methods often rely on influence-estimation heuristics: Engstrom et al. (2024) leveraged linear datamodel scores (Park et al., 2023) to select data for LLMs pretraining; Xia et al. leveraged influence estimation to select data for LLMs fine tuning. Hu et al. (2024) highlighted these methods’ limitations and proposed iterative heuristics. A concurrent work (Liu et al., ) also frames data selection as an optimal transport problem. However it still follows the same gradient distance estimation in (Xia et al., ) and uses a continuous OT-based formulation, making it less efficient and effective. Our approach introduces a novel distance estimation method and a greedy subset selection pipeline to address these challenges.
@@ Acknowledgments
Impact Statement
This paper presents TAROT, a targeted data selection framework using Optimal Transport to improve model efficiency and performance across diverse tasks. By addressing biases in influence-based heuristics, TAROT enables better dataset alignment with target distributions, reducing computational costs and potential environmental impact. This approach can enhance data efficiency, benefiting scenarios where data privacy or accessibility is a concern. However, like any selection strategy, it may introduce biases if not carefully managed. Future research should explore its implications on fairness and generalization, but overall, TAROT represents a step toward more efficient and responsible data selection in machine learning.
Appendix A More Related Work
Data Influence Estimation. Pruthi et al. (2020) proposed a method to estimate data influence by tracing the gradient descent process. Another approach utilizes datamodels: Ilyas et al. (2022) showed that simple linear datamodels can predict model outputs effectively. Building on this, Park et al. (2023) introduced a scalable technique for estimating datamodel scores. Subsequent works have optimized these methods, focusing on checkpoint ensembles (Deng et al., 2024), efficient gradient storage (Zhang et al., 2024), and scalable gradient projection (Choe et al., 2024). Our method extends Pruthi et al. (2020) by enhancing estimation robustness through feature covariance whitening.
Optimal Transport for Data Selection. Optimal transport has been used for dataset similarity estimation (Alvarez-Melis & Fusi, 2020) and sensitivity analysis in data valuation (Just et al., 2023). Kang et al. (2024) applied it to optimize data selection ratios across multiple sources. We integrate influence estimation with optimal transport, enabling effective subset selection to boost target domain performance.
Appendix B More Implementation Details
B.1 Influence Estimation
ZCA Whitening. As discussed in Section 4.1, we use ZCA whitening to achieve higher LDS scores due to its ability to preserve the data structure. However, our empirical analysis revealed that ZCA whitening has minimal impact on data selection performance. Therefore, for the data selection experiments, we opted for the more computationally efficient Cholesky whitening.
@@ B.1 Influence Estimation
Linear Data Modeling Score (LDS). As mentioned in Sec. 4.1, we employ LDS as evaluation metric for influence estimation. LDS is computed by the Spearman rank correlation between the model outputs of models trained on randomly selected subsets and the average of attribution scores as predictions of those model outputs. For Cifar-10, we directly reuse the provided outputs of models trained on different subsets in the Github repository of TRAK (Park et al., 2023), and utilize these outputs to calculate LDS. For nuScences, we randomly select 100 subsets of the candidate datasets. Each subset is sampled to be 50% of the size of the candidate dataset. Then we independently train 100 models on these subsets and utilize their model outputs to calculate LDS.
@@
B.2 Semantic Segmentation
Table 3 shows detailed hyperparameters for the semantic segmentation experiments.
B.3 Motion Prediction
We use the UniTraj (Feng et al., 2024) framework for motion prediction experiments. Detailed experiment settings are provided in Table 4.
B.4 Instruction Tuning
We follow the settings and hyperparameters in LESS (Xia et al., ). Details are provided in Table 5.
Appendix C Details for Algorithms
We present the detailed data selection algorithm described in Section 3.3 in Algorithms 1 and 2
@@ Appendix C Details for Algorithms
Table 6: Performance comparison (↑ means higher is better). All methods use 5% of the full training data. TAROT-OTM shows relative improvement over TAROT-Fixed in parentheses.
Dataset LLaMA-3.1-8B MMLU LLaMA-3.1-8B BBH Qwen-2.5-7B MMLU Qwen-2.5-7B BBH
5% LESS 65.7 62.6 74.3 66.3
5% TSDS 65.2 63.1 73.9 66.2
5% TAROT 66.0 65.0 74.1 66.9
TAROT-OTM 65.7 (0.13%) 63.6 (0.21%) 74.3 (0.09%) 68.9 (0.13%)
Method Ranking ($\downarrow$) minADE5555 ($\downarrow$)
Wayformer-TAROT 1 0.92
Wayformer-All (Nayakanti et al., 2023) 2 0.94
MTR-UniTraj (Feng et al., 2024) 3 0.96
Goal-LBP (Yao et al., 2023) 4 1.02
SemanticFormer (Sun et al., 2024) 5 1.14
Table 8: The wall clock runtime (measured as single H100 GPU hours) of TAROT compared with LESS and TSDS on instruction tuning task. The gradient feature computation stage is the same for all the methods.
Gradient Features Computation Data Selection
LESS 32 Hours 46 Seconds
TAROT-Fixed 59 Seconds
TAROT-OTM 118 Seconds
TSDS 10 Hours
Table 9: Asymptotic complexity analysis.
Gradient Features Computation Fixed-Size/OTM Selection
Complexity $\mathcal{O}(|\mathcal{D}|\cdot N)$ $\mathcal{O}\left(|\mathcal{D}|\cdot\left|\mathcal{D}_{\text{tar }}\right|\cdot d\right)$
Appendix D NuScenes Leaderboard Performance
In the motion prediction task, TAROT helps Wayformer achieve the first place on the nuScenes leaderboard (Table 7). surpassing MTR-UniTraj (Feng et al., 2024), which relies on large-scale datasets ($2000k$ samples). These results demonstrate that targeted data selection offers a promising alternative to data scaling for breaking performance boundaries.
Appendix E Discussion and Comparison with TSDS
TSDS frames data selection as an optimal transport (OT) problem, incorporating a diversity regularizer via kernel density estimation. It selects samples based on distances in gradient embedding space.
In contrast, TAROT uses Whitened Feature Distance (WFD), which removes the confounding effects of gradient covariance and scale. This provides more robust and accurate feature distance estimates, leading to better influence estimation—surpassing the state-of-the-art TRAK method. Besides, while TSDS uses a continuous OT-based formulation and supports subset size control via sampling, TAROT differs in its greedy, deterministic subset selection that explicitly minimizes the empirical OT distance at each iteration. This enables stronger control over the selected subset and much faster speed for data selection. Indeed, we empirically found that TSDS costs significantly more time than TAROT. The wall-clock time analysis is shown in Table 8. Finally, TAROT introduces an early stopping criterion for OT-based selection via tracking OT distance increase, allowing estimation of optimal selection ratios, which TSDS does not address.
We also include experimental comparisons with TSDS, following the same evaluation settings as described in Section 4.4. The results are shown in Table 6.
Appendix F More Qualitative Results of WFD
@@
We provide additional qualitative results of WFD in Figures 9 and 10. One clear observation is that the most supportive samples tend to closely resemble the target sample in both the input and output spaces. In motion prediction, the most negative samples also share similarities with the target sample in the input space (e.g., map and history trajectories); however, their ground truth trajectories often exhibit minor differences. For instance, in the first row of samples, the most negative examples demonstrate behaviors such as turning left or driving in the right-most lanes, whereas the target sample follows a trajectory in the lane adjacent to the right-most lane. These qualitative findings align with the definition of influence estimation: the most positive samples guide the model to make predictions similar to the target sample, while the negative samples encourage divergent predictions.
Appendix G Computational Complexity Analysis
Table 9 presents the asymptotic complexity of TAROT. The most computationally expensive step is the computation of gradient features, with its cost scaling linearly with the candidate dataset size $|\mathcal{D}|$, the number of checkpoints $N$, and the gradient dimension $d$. However, the computational costs incurred during these two stages are one-time expenses that can be amortized across multiple target tasks. The data selection process requires minimal computation.
Table 8 demonstrates the wall-clock runtime (in single H100 GPU hours) of TAROT compared to LESS and TSDS for the instruction tuning task. The gradient feature computation time is identical for both methods. For data selection, both TAROT-Fixed and LESS complete in under one minute, while TAROT-OTM incurs an additional runtime of approximately one minute.